Can someone help me with these two problems?
(1)
Add. Write the answer in the lowest terms.
3/8 + 5/8
(2)
Add. Write the answer in lowest terms.
5/9 + 8/9
Thank you.

Answer by fractalier(2101)

If the denominators are the same, you merely add numerators...so we get
1) 3/8 + 5/8 = 8/8 = 1
2) 5/9 + 8/9 = 13/9 = 1 4/9